what is value of cosec(A+B) in terms of cosec and sec

Grade:10

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

Cosec(A+B)
=1/ sin(A+B)
=1/(sin A cos B + cosA sinB)
=1/((cosecA)(secB)+(1/secA)(1/cosecB))

Now you can take LCM in the denomenator and solve it